# Radiative Cooling Effects on Plasmoid Formation in Black Hole Accretion Flows with Multiple Magnetic Loops

### v1

@@ Abstract
Results. Our results demonstrate that radiative cooling suppresses the transition to the magnetically arrested disk (MAD) state by reducing magnetic flux accumulation near the horizon. It significantly modifies the disk morphology by lowering the electron temperature and compressing the disk, which leads to increased density at the equatorial plane and decreased magnetization. Within the current sheets, radiative cooling triggers layer compression and the collapse of plasmoids, shortening their lifetime and reducing their size, while the frequency of plasmoid events increases. Moreover, we observe enhanced negative energy-at-infinity density in plasmoids near the ergosphere, with its peaks corresponding to plasmoid presence.
@@ 1 Introduction
Despite extensive efforts, the precise mechanism responsible for these flares remains under debate. Several mechanisms have been proposed, including flux eruption events (Porth et al. 2021; Antonopoulou et al. 2025), magnetic reconnection (Lin and Yuan 2024; Lin et al. 2023), and plasmoid chain along current sheet (Dimitropoulos et al. 2025). These processes collectively contribute to the observed variability and energetic phenomena near the BH (Trippe et al. 2007; Yuan et al. 2004). Observations from the Event Horizon Telescope (EHT) on Sgr A suggest the presence of a strong and ordered magnetic field (Event Horizon Telescope Collaboration et al. 2024). These results indicate a preference for the magnetically arrested disk (MAD) regime compared to the Standard and Normal Evolution (SANE) regime and other types of accretion flows (Event Horizon Telescope Collaboration et al. 2022a, 2024, b). The absence of a powerful jet in current observations, typically associated with the MAD accretion regime (Chavez et al. 2024; Narayan et al. 2022), suggests the incompleteness of our current models.
@@
Large-scale magnetized jets generated by accretion flows around BHs often exhibit variable energy output (Nathanail et al. 2020). Due to the inherently turbulent nature of the accretion process (Liska et al. 2022), magnetic field loops of different polarities are generated around the BH. The merge of the magnetic loops causes frequent polarity inversion events and forms current sheets (Ripperda et al. 2022; El Mellah et al. 2023; Kumar et al. 2015). The tearing instability in the current sheets creates plasmoids and x-points, making them ideal places for particle acceleration and the flaring events of Sgr A via magnetic reconnection(Tchekhovskoy et al. 2009; Ripperda et al. 2017, 2019; GRAVITY Collaboration et al. 2018; Gravity Collaboration et al. 2023; Ripperda et al. 2022).
@@
High-resolution GRMHD simulations show that mean-field dynamo and magnetorotational instability can transform toroidal magnetic fields into poloidal loops, creating a multi-loop magnetic configuration in accretion tori (Liska et al. 2020; Kaufman et al. 2023; Gottlieb et al. 2023; Rodman and Reynolds 2024). To mimic this process, simulations with initial multiloop magnetic field configuration have been performed to study the dynamics of the accretion flows (Nathanail et al. 2020, 2022; Jiang et al. 2023, 2024). The alternating polarity configuration not only prevents the accumulation of magnetic fields but also facilitates frequent reconnection between magnetic loops of opposite polarities, thereby releasing magnetic energy. A series of two-dimensional (2D) GRMHD simulations with multi-loop configurations was conducted by Nathanail et al. (2020), while three-dimensional (3D) simulations with multi-loop magnetic field configurations were performed by Chashkina et al. (2021) and Nathanail et al. (2022). These studies demonstrate that the multi-loop magnetic field configuration can efficiently produce energetic plasmoids via magnetic reconnection, which have been proposed to explain flaring events from Sgr A (Petropoulou et al. 2016; Zhang et al. 2018; Christie et al. 2019). Another possible explanation for these flares is the energy release during flux eruption events in MAD state, which also provide suitable conditions for accelerating non-thermal particles (Chashkina et al. 2021; Scepi et al. 2022; Jiang et al. 2023; Grigorian and Dexter 2024). Different with multi-loop models which is usually in the SANE state with a weak jet, MAD models always launches powerful jets. However, strong observational evidence for such a strong jet in Sgr A remains unclear (Event Horizon Telescope Collaboration et al. 2022a).
@@
With the advancement of astronomical observations and numerical simulations, BHs have emerged as unique laboratories for testing plasma electrodynamics under extreme conditions (Lasota et al. 2014). Among the most profound predictions of general relativity is the possibility of extracting rotational energy from a BH via the inflow of negative energy and angular momentum across the event horizon (Penrose and Floyd 1971). Recently, the nonlinear nature of magnetic reconnection and plasma dynamics near the surface of the accretion tori around BHs has been increasingly recognized (Asenjo and Comisso 2015), and it has been reported that energy extraction can happen in the formation of plasmoid chains via Penrose process (PP) (Fan et al. 2024; Shen and YuChih 2025). Most of these studies have been conducted within purely analytical frameworks. In this work, we extend the investigation by performing numerical simulations and analysis within a multi-loop configuration that incorporates the effects of radiative cooling.
@@ 2 Numerical setup
Our GRMHD simulations are conducted with the BHAC code (Porth et al. 2017; Olivares et al. 2019). BHAC code solves the ideal MHD equations in geometric units ($G=M=c=1$), which is given in equations expressed in covariant notation as follows:
@@ 2.1 Model setup
Our simulations use spherical modified Kerr–Schild coordinate. The simulation domain extends radially from the BH event horizon to $r=2500\,r_{\rm g}$ with a logarithmic spacing. In the polar direction, the domain spans from the polar angle $\theta$ = 0 to $\pi$ with uniform spacing. In all 2D simulations, we adopt a standard resolution of $1024\times 512$, except for one high-resolution case with an effective resolution of $4096\times 2048$ (see Appendix B for details on the numerical setup). For the 3D simulations, we employ two levels of static mesh refinement, achieving an effective resolution of $320\times 160\times 160$. The 2D simulations are evolved until $t=15\,000\,\rm M$, while the 3D simulations are extended to $t=25\,000\,\rm M$ to reach quasi-steady state.
@@ 2.2 Radiative cooling
Following Sądowski et al. (2017); Ressler et al. (2015), we solve the electron-entropy equation in the presence of dissipative heating, Coulomb interaction, and radiative cooling processes:
@@
3 Influence of Radiative-cooling to Accretion Flows: RESULTS
@@ 3.1 Comparison of properties in the accretion flow
In Fig. 1, the time evolutions of mass accretion rate $\dot{M}$ and normalized magnetic flux $\Phi_{\rm BH}/\sqrt{\dot{M}}$ of the models are presented. Panels (a) and (c) of Fig. 1 depict models initialized with smaller magnetic loops (wavelength $\lambda_{r}=40$), while panels (b) and (d) show models with larger magnetic loops (wavelength $\lambda_{r}=80$). Different colors in these panels indicate variations in the values of mass accretion rate normalized by Eddington ratio, $\dot{M}_{\rm Edd}$, and with/without radiative cooling. The dash line in the panels represent $\Phi_{\text{BH}}/\sqrt{\dot{M}}=15$, which is the criterion for MAD regime (Tchekhovskoy et al. 2011; Mizuno et al. 2021).
@@
From the magnetic flux evolutions in Fig. 1, we see a decreased magnetic flux from the models with stronger cooling. The vertical blue lines in the panel indicates the moment the accretion flow achieves a quasi-steady state for smaller magnetic loop cases. The horizontal lines represent the average normalized magnetic flux during this period. In the quasi-steady state, the non-cooling case (red line) exhibits greater magnetic flux accumulation compared to the case with radiative cooling (blue line). Due to the relatively smaller magnetic loops, the accretion flows do not reach the MAD regime (Jiang et al. 2023). Radiative cooling mostly reduces the magnetic flux, which is also evident in models D, E, and F in the larger loop runs in Fig. 1(d). We note, however, that the cooling-induced reduction of the normalized magnetic flux at the event horizon may not be universal. Recent work by Singh et al. (2025) demonstrated that the response of the MAD parameter to radiative cooling can depend sensitively on the black hole spin, with different spin values exhibiting distinct trends in magnetic flux accumulation. In the present study, we restrict our analysis to a single black hole spin configuration and therefore do not explore this additional parameter space. We focus on a rapidly spinning black hole with $a=0.9375$, motivated by observational constraints suggesting that Sgr A likely possesses a high spin (Event Horizon Telescope Collaboration et al. 2022b). Consequently, while our results indicate a systematic suppression of magnetic flux with increasing cooling strength for the spin considered here, a broader survey over black hole spins will be required to establish the generality of this behavior.
@@
For the larger loop models (wavelength $\lambda_{r}=80\,r_{\rm g}$), shown in the right panels of Fig. 1, fewer magnetic loops within the torus result in reduced magnetic dissipation between them. The merging of these loops generates abundant small-scale magnetic fields, which initially prevent significant magnetic field accumulation near the BH, producing a SANE-type flow in the early stages of the simulations. As these small-scale fields coalesce into a more ordered, large-scale magnetic structure, the accretion flow transitions from a SANE to a MAD regime, as observed in Panel (d) of Fig. 1 after the vertical lines. The light-blue and light-red vertical lines indicate the transition time from SANE to MAD regime. From model E (light-blue lines), which has a relatively weak cooling effect, we see a delayed transition to the MAD regime compared to the non-cooling case (model F, light-red lines). While in model D (green lines), the cooling effect is so strong that the SANE to MAD transition does not appear in the simulation range. Radiative cooling reduces the accumulation of magnetic flux near the BH horizon, resulting in insufficient magnetic flux pile-up to reach the MAD regime.
@@ 3.2.1 Electron temperature
Figure 3: Time-averaged electron temperature $\Theta_{\,\rm e}$, plasma $\beta$, and density $\rho$ distributions for three models. From top to bottom, the panels correspond to models A, B, and C. The plasma beta parameter is defined as $\beta=p_{\,\rm g}/p_{\text{mag}}$, where $p_{\,\rm g}$ represents the gas pressure, and $p_{\text{mag}}=b^{2}/2$ denotes the magnetic pressure. The light black contours in the density profile represent the magnetic field. The average is taken over the time range $t=8000\,\rm M$ to $13000\,\rm M$.
@@
Fig. 5 presents azimuthal- and time-averaged polar-angle profiles of the ion-to-electron temperature ratio of different cooling models at $r=15\,\rm M$. We observe that both model A (with cooling) and model C (without cooling) exhibit a higher ion-to-electron temperature ratio near the equatorial plane compared to other polar angles; similar plots have been shown by Mizuno et al. (2021). Model A tends to exhibit a globally higher ion-to-electron temperature ratio compared to both the no-cooling case (model C) and the weaker cooling case (model B). It is worth noting that in both model B and model C, the ion temperature remains lower than the electron temperature ($T_{\,\rm i}<T_{\,\rm e}$) throughout the domain. In contrast, model A exhibits a region near the equatorial plane where $T_{\,\rm i}>T_{\,\rm e}$, due to the radiative cooling, which primarily reduces the electron temperature in the disk region. While model A still showing $T_{\,\rm i}<T_{\,\rm e}$ toward the polar regions, consistent with the other models. This behavior arises because most of the electron heating occurs around the funnel region. Consistent with our expectations, the high-resolution simulations reproduce similar trends observed in the low-resolution runs, confirming the influence of radiative cooling on the thermodynamic structure of the accretion flows.
@@ 3.2.2 Radiative collapse and plasmoid formation
Figure 8: Time evolution of a plasmoid chain in high-resolution model hr over the time interval $t=8170\,\rm M$ to $t=8180\,\rm M$. The upper panel shows electron temperature, and the lower panel presents density. We mark a red box for the place of the plasmoid chain in each panel.
@@
As radiative losses deplete internal energy from the reconnection layer, its temperature decreases. In response, the density increases to maintain pressure equilibrium, as previously discussed, ultimately leading to collapse. In addition to radiative collapse, the reconnection layer may also be possible to grow local radiative cooling instabilities (Field 1965). These instabilities arise from pressure perturbations that become unstable under radiative cooling. The interaction between these thermal and the tearing instabilities can play a significant role in shaping the transient dynamics of the reconnection process (Forbes and Malherbe 1991). In our simulations, we observe a collapse induced by radiative cooling.
@@ 3.2.3 Statistical properties of plasmoid
In order to analyze the statistical properties of plasmoids with radiative cooling, we try to detect them in the simulations. We apply a plasmoid detection method based on Nathanail et al. (2020), by tracking blob-like structures in our simulations using the scikit-image package (Van der Walt et al. 2014). This approach allows us to capture the positions and sizes of the plasmoids within a selected region, which extends from 0 to $20\,r_{\rm g}$ in $x$-direction and from 20-20 to $20\,r_{\rm g}$ in $y$-direction of the snapshots with a cadence of $10\,\rm M$. We applied the same detection criteria to all models to ensure generalizability. We also constrain the regions where the Bernoulli constant ($hu_{t}$) is greater than 1.02 for this plasmoid detection.
As discussed in Sect. 3.2.2, radiative cooling leads to the collapse of the plasmoids. This implies that the amount of detected plasmoids decreases after the collapse occurs. In Fig. 10, we plot the time evolution of number of plasmoids in different models. In the presence of radiative cooling, radiative collapse causes the plasmoid to dissipate and gradually vanish. Compared to cases without radiative cooling, this collapse shortens the lifetime of the plasmoid chain. This results in a more spiky structure in the plasmoid count evolution.
@@
To facilitate a more quantitative discussion, we perform a statistical analysis of the transient behaviors of the plasmoid chains, which is shown in Fig. 11. Here, we define the duration based on the criterion of the mean value minus standard deviation. The time points where the count crosses below this threshold are identified as the start and end of individual plasmoid events. This result aligns with our expectations: radiative collapse leads to a shorter lifetime of the plasmoid chains. Additionally, the increased density and compressed layer due to radiative cooling further increase the magnetic reconnection rate, yielding more frequent production of plasmoid chains. We note that the average number of plasmoids is not influenced significantly by the cooling effect, which we attribute to the combined effect of these two processes. To complement the temporal diagnostics, we conduct a Fourier analysis over the same time interval to investigate variability associated with plasmoid formation, as presented in Fig. 12 and summarize the mean amplitudes in three different frequencies in Table 2.
@@
To further explore the impact of radiative cooling on the plasmoids, we analyze and compare the radii of plasmoids detected in models A, B, and C, shown in Fig. 13. The results reveal that the layer compression caused by radiative cooling also affects plasmoids, resulting in tendency for more production of smaller plasmoids. The mean radius of Models A, B, and C is $r=1.43$, $1.59$, and $1.80\,\rm G\rm M/\rm c^{2}$. It is reflected in the increase of the total number of plasmoids seen in Fig. 10.

### v2

@@ Abstract
Results. Our results demonstrate that radiative cooling suppresses the transition to the magnetically arrested disk (MAD) state by reducing magnetic flux accumulation near the horizon. It significantly modifies the disk morphology by lowering the electron temperature and compressing the disk, which leads to increased density at the equatorial plane and decreased magnetization. Within the current sheets, radiative cooling triggers layer compression and the collapse of plasmoids, shortening their lifetime and reducing their size, while the frequency of plasmoid events increases. Moreover, we observe enhanced negative energy-at-infinity density in plasmoids near the ergosphere, with its peaks corresponding to plasmoid formation events.
@@ 1 Introduction
Despite extensive efforts, the precise mechanism responsible for these flares remains under debate. Several mechanisms have been proposed, including flux eruption events (Porth et al. 2021; Antonopoulou et al. 2025), magnetic reconnection (Lin and Yuan 2024; Lin et al. 2023), and plasmoid chains along current sheet (Dimitropoulos et al. 2025). These processes collectively contribute to the observed variability and energetic phenomena near the BH (Trippe et al. 2007; Yuan et al. 2004). Observations from the Event Horizon Telescope (EHT) on Sgr A suggest the presence of a strong and ordered magnetic field (Event Horizon Telescope Collaboration et al. 2024). These results indicate a preference for the magnetically arrested disk (MAD) regime compared to the Standard and Normal Evolution (SANE) regime and other types of accretion flows (Event Horizon Telescope Collaboration et al. 2022a, 2024, b). The absence of a powerful jet in current observations, typically associated with the MAD accretion regime (Chavez et al. 2024; Narayan et al. 2022), suggests the incompleteness of our current models.
@@
Large-scale magnetized jets generated by accretion flows around BHs often exhibit variable energy output (Nathanail et al. 2020). Due to the inherently turbulent nature of the accretion process (Liska et al. 2022), magnetic field loops of different polarities are generated around the BH. The merging of the magnetic loops causes frequent polarity inversion events and forms current sheets (Ripperda et al. 2022; El Mellah et al. 2023; Kumar et al. 2015). The tearing instability in the current sheets creates plasmoids and x-points, making them ideal places for particle acceleration and the flaring events of Sgr A via magnetic reconnection(Tchekhovskoy et al. 2009; Ripperda et al. 2017, 2019; GRAVITY Collaboration et al. 2018; Gravity Collaboration et al. 2023; Ripperda et al. 2022).
@@
High-resolution GRMHD simulations show that mean-field dynamo and magnetorotational instability can transform toroidal magnetic fields into poloidal loops, creating a multi-loop magnetic configuration in accretion tori (Liska et al. 2020; Kaufman et al. 2023; Gottlieb et al. 2023; Rodman and Reynolds 2024). To mimic this process, simulations with initial multiloop magnetic field configuration have been performed to study the dynamics of the accretion flows (Nathanail et al. 2020, 2022; Jiang et al. 2023, 2024). The alternating polarity configuration not only prevents the accumulation of magnetic fields but also facilitates frequent reconnection between magnetic loops of opposite polarities, thereby releasing magnetic energy. A series of two-dimensional (2D) GRMHD simulations with multi-loop configurations was conducted by Nathanail et al. (2020), while three-dimensional (3D) simulations with multi-loop magnetic field configurations were performed by Chashkina et al. (2021) and Nathanail et al. (2022). These studies demonstrate that the multi-loop magnetic field configuration can efficiently produce energetic plasmoids via magnetic reconnection, which have been proposed to explain flaring events from Sgr A (Petropoulou et al. 2016; Zhang et al. 2018; Christie et al. 2019). Another possible explanation for these flares is the energy release during flux eruption events in MAD state, which also provide suitable conditions for accelerating non-thermal particles (Chashkina et al. 2021; Scepi et al. 2022; Jiang et al. 2023; Grigorian and Dexter 2024). Unlike SANE models, which typically produce only weak jets, MAD models generally launch powerful jets. However, strong observational evidence for such a powerful jet in Sgr A remains unclear (Event Horizon Telescope Collaboration et al. 2022a).
@@
With the advancement of astronomical observations and numerical simulations, BHs have emerged as unique laboratories for testing plasma electrodynamics under extreme conditions (Lasota et al. 2014). Among the most profound predictions of general relativity is the possibility of extracting rotational energy from a BH via the inflow of negative energy and angular momentum across the event horizon (Penrose and Floyd 1971). Recently, the nonlinear nature of magnetic reconnection and plasma dynamics near the surface of the accretion tori around BHs has been increasingly recognized (Asenjo and Comisso 2015), and it has been reported that energy extraction can occur in the formation of plasmoid chains via Penrose process (PP) (Fan et al. 2024; Shen and YuChih 2025). Most of these studies have been conducted within purely analytical frameworks. In this work, we extend the investigation by performing numerical simulations and analysis within a multi-loop configuration that incorporates the effects of radiative cooling.
@@ 2 Numerical setup
Our GRMHD simulations are conducted with the BHAC code (Porth et al. 2017; Olivares et al. 2019). BHAC code solves the ideal MHD equations in geometric units ($G=M=c=1$), which are given in equations expressed in covariant notation as follows:
@@ 2.1 Model setup
Our simulations use spherical modified Kerr–Schild coordinates. The simulation domain extends radially from the BH event horizon to $r=2500\,r_{\rm g}$ with a logarithmic spacing. In the polar direction, the domain spans from the polar angle $\theta$ = 0 to $\pi$ with uniform spacing. In all 2D simulations, we adopt a standard resolution of $1024\times 512$, except for one high-resolution case with an effective resolution of $4096\times 2048$ (see Appendix B for details on the numerical setup). For the 3D simulations, we employ two levels of static mesh refinement, achieving an effective resolution of $320\times 160\times 160$. The 2D simulations are evolved until $t=15\,000\,\rm M$, while the 3D simulations are extended to $t=25\,000\,\rm M$ to reach quasi-steady state.
@@ 2.2 Radiative cooling
Following Ressler et al. (2015) and Sądowski et al. (2017), we solve the electron-entropy equation in the presence of dissipative heating, Coulomb interaction, and radiative cooling processes:
@@
3 Influence of radiative cooling on accretion flows
@@ 3.1 Comparison of properties in the accretion flow
In Fig. 1, the time evolutions of mass accretion rate $\dot{M}$ and normalized magnetic flux $\Phi_{\rm BH}/\sqrt{\dot{M}}$ of the models are presented. Panels (a) and (c) of Fig. 1 depict models initialized with smaller magnetic loops (wavelength $\lambda_{r}=40$), while panels (b) and (d) show models with larger magnetic loops (wavelength $\lambda_{r}=80$). Different colors in these panels indicate variations in the values of mass accretion rate normalized by Eddington ratio, $\dot{M}_{\rm Edd}$, and with/without radiative cooling. The dashed lines in the panels represent $\Phi_{\text{BH}}/\sqrt{\dot{M}}=15$, which is the criterion for MAD regime (Tchekhovskoy et al. 2011; Mizuno et al. 2021).
@@
From the magnetic flux evolutions in Fig. 1, we observe a decreased magnetic flux from the models with stronger cooling. The vertical blue lines in the panel indicates the moment the accretion flow achieves a quasi-steady state for smaller magnetic loop cases. The horizontal lines represent the average normalized magnetic flux during this period. In the quasi-steady state, the non-cooling case (red line) exhibits greater magnetic flux accumulation compared to the case with radiative cooling (blue line). Due to the relatively smaller magnetic loops, the accretion flows do not reach the MAD regime (Jiang et al. 2023). Radiative cooling mostly reduces the magnetic flux, which is also evident in models D, E, and F in the larger loop runs in Fig. 1(d). We note, however, that the cooling-induced reduction of the normalized magnetic flux at the event horizon may not be universal. Recent work by Singh et al. (2025) demonstrated that the response of the MAD parameter to radiative cooling can depend sensitively on the black hole spin, with different spin values exhibiting distinct trends in magnetic flux accumulation. In the present study, we restrict our analysis to a single black hole spin configuration and therefore do not explore this additional parameter space. We focus on a rapidly spinning black hole with $a=0.9375$, motivated by observational constraints suggesting that Sgr A likely possesses a high spin (Event Horizon Telescope Collaboration et al. 2022b). Consequently, while our results indicate a systematic suppression of magnetic flux with increasing cooling strength for the spin considered here, a broader survey over black hole spins will be required to establish the generality of this behavior.
@@
For the larger loop models (wavelength $\lambda_{r}=80\,r_{\rm g}$), shown in the right panels of Fig. 1, fewer magnetic loops within the torus result in reduced magnetic dissipation between them. The merging of these loops generates abundant small-scale magnetic fields, which initially prevent significant magnetic field accumulation near the BH, producing a SANE-type flow in the early stages of the simulations. As these small-scale fields coalesce into a more ordered, large-scale magnetic structure, the accretion flow transitions from a SANE to a MAD regime, as observed in Panel (d) of Fig. 1 after the vertical lines. The light-blue and light-red vertical lines indicate the transition time from SANE to MAD regime. From model E (light-blue lines), which has a relatively weak cooling effect, we see a delayed transition to the MAD regime compared to the non-cooling case (model F, light-red lines). While in model D (green lines), the cooling effect is so strong that the SANE to MAD transition does not occur within the simulation duration. Radiative cooling reduces the accumulation of magnetic flux near the BH horizon, resulting in insufficient magnetic flux pile-up to reach the MAD regime.
@@ 3.2.1 Electron temperature
Figure 3: Time-averaged electron temperature $\Theta_{\rm e}$, plasma $\beta$, and density $\rho$ distributions for three models. From top to bottom, the panels correspond to models A, B, and C. The plasma beta parameter is defined as $\beta=p_{\rm g}/p_{\rm mag}$, where $p_{\rm g}$ represents the gas pressure, and $p_{\rm mag}=b^{2}/2$ denotes the magnetic pressure. The light black contours in the density profile represent the magnetic field. The average is taken over the time range $t=8000\,{\rm M}$ to $13000\,{\rm M}$.
@@
Fig. 5 presents azimuthal- and time-averaged polar-angle profiles of the ion-to-electron temperature ratio of different cooling models at $r=15\,\rm M$. We observe that both model A (with cooling) and model C (without cooling) exhibit a higher ion-to-electron temperature ratio near the equatorial plane compared to other polar angles; similar plots have been shown by Mizuno et al. (2021). Model A tends to exhibit a globally higher ion-to-electron temperature ratio compared to both the no-cooling case (model C) and the weaker cooling case (model B). It is worth noting that in both model B and model C, the ion temperature remains lower than the electron temperature ($T_{\,\rm i}<T_{\,\rm e}$) throughout the domain. In contrast, model A exhibits a region near the equatorial plane where $T_{\,\rm i}>T_{\,\rm e}$, due to the radiative cooling, which primarily reduces the electron temperature in the disk region. While model A still shows $T_{\,\rm i}<T_{\,\rm e}$ toward the polar regions, consistent with the other models. This behavior arises because most of the electron heating occurs around the funnel region. Consistent with our expectations, the high-resolution simulations reproduce similar trends observed in the low-resolution runs, confirming the influence of radiative cooling on the thermodynamic structure of the accretion flows.
@@ 3.2.2 Radiative collapse and plasmoid formation
Figure 8: Time evolution of a plasmoid chain in high-resolution model hr over the time interval $t=8170\,\rm M$ to $t=8180\,\rm M$. The upper panel shows electron temperature, and the lower panel presents density. We mark the location of the plasmoid chain with a red box in each panel.
@@
As radiative losses deplete internal energy from the reconnection layer, its temperature decreases. In response, the density increases to maintain pressure equilibrium, as previously discussed, ultimately leading to collapse. In addition to radiative collapse, the reconnection layer may also be subject to local radiative cooling instabilities (Field 1965). These instabilities arise from pressure perturbations that become unstable under radiative cooling. The interaction between these thermal and the tearing instabilities can play a significant role in shaping the transient dynamics of the reconnection process (Forbes and Malherbe 1991). In our simulations, we observe a collapse induced by radiative cooling.
@@ 3.2.3 Statistical properties of plasmoid
In order to analyze the statistical properties of plasmoids with radiative cooling, we aim to detect them in the simulations. We apply a plasmoid detection method based on Nathanail et al. (2020), by tracking blob-like structures in our simulations using the scikit-image package (Van der Walt et al. 2014). This approach allows us to capture the positions and sizes of the plasmoids within a selected region, which extends from 0 to $20\,r_{\rm g}$ in $x$-direction and from 20-20 to $20\,r_{\rm g}$ in $y$-direction of the snapshots with a cadence of $10\,\rm M$. We applied the same detection criteria to all models to ensure generalizability. We also constrain the regions where the Bernoulli constant ($hu_{t}$) is greater than 1.02 for this plasmoid detection.
As discussed in Sect. 3.2.2, radiative cooling leads to the collapse of the plasmoids. This implies that the number of detected plasmoids decreases after the collapse occurs. In Fig. 10, we plot the time evolution of number of plasmoids in different models. In the presence of radiative cooling, radiative collapse causes the plasmoid to dissipate and gradually vanish. Compared to cases without radiative cooling, this collapse shortens the lifetime of the plasmoid chain. This results in a more spiky structure in the plasmoid count evolution.
@@
To facilitate a more quantitative discussion, we perform a statistical analysis of the transient behaviors of the plasmoid chains, which is shown in Fig. 11. Here, we define the duration based on the criterion of the mean value minus standard deviation. The time points where the count crosses below this threshold are identified as the start and end of individual plasmoid events. This result aligns with our expectations: radiative collapse shortens the lifetime of the plasmoid chains. Additionally, the increased density and compressed layer due to radiative cooling further increase the magnetic reconnection rate, yielding more frequent production of plasmoid chains. We note that the average number of plasmoids is not influenced significantly by the cooling effect, which we attribute to the combined effect of these two processes. To complement the temporal diagnostics, we conduct a Fourier analysis over the same time interval to investigate variability associated with plasmoid formation, as presented in Fig. 12 and summarize the mean amplitudes in three different frequencies in Table 2.
@@
To further explore the impact of radiative cooling on the plasmoids, we analyze and compare the radii of plasmoids detected in models A, B, and C, shown in Fig. 13. The results reveal that the layer compression caused by radiative cooling also affects plasmoids, resulting in a tendency toward the production of smaller plasmoids. The mean radius of Models A, B, and C is $r=1.43$, $1.59$, and $1.80\,\rm G\rm M/\rm c^{2}$. It is reflected in the increase of the total number of plasmoids seen in Fig. 10.